OpenStudy (anonymous):

Evaluate: 32 + (8 − 2) • 4 − 6 over 3

OpenStudy (johnweldon1993):

So does this look like \[\frac{ 32 + (8 - 2) \times 4 - 6}{ 3 }\] or \[32 + (8 - 2) \times 4 - \frac{ 6 }{ 3 }\] ??

OpenStudy (anonymous):

its 6/3

OpenStudy (anonymous):

so the second one

OpenStudy (dan815):

johnweldon!

OpenStudy (dan815):

do u know how to play checkers

OpenStudy (johnweldon1993):

Okay...so we follow the order of operations here PEMDAS so do what is in the parenthesis first (8 - 2) = 6 right...so we have 32 + 6 x 4 - 6/3 Next we do multiplication and then division 6 x 4 = 24 6 / 3 = 2 so 32 + 24 - 2 so what is 32 + 24 - 2?

OpenStudy (johnweldon1993):

And yes I do @dan815 lol *not like pro or anything but yeah haha*

OpenStudy (dan815):

just asking cuz i randomly found a site where u can play checkers and chess today without having to sign up or anything so looking to see if anyone wanna play lol

OpenStudy (dan815):

i wanna play a game of monopoly with ppl i know on here lol that will be fun, gotta find a site that lets u play monopoly without signing up now

OpenStudy (anonymous):

54?

OpenStudy (johnweldon1993):

Looks correct to me!